OpenStudy (anonymous):

true or false: x^2-16x+64/x-8

OpenStudy (solomonzelman):

HINT: x²-16x+64 = (x-8)²

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \frac{(x-8)^2}{x-8}=?}\)

OpenStudy (solomonzelman):

This is not a true or false question though, so I can't do anything

OpenStudy (anonymous):

does it have a remainder

OpenStudy (anonymous):

No, it doesn't. (x-8)^2=x^2-16x+64 exactly.